OpenStudy (anonymous):

: Write an expression for the following real-world situation. If Mr. Jackson earns $1000 for every used car he sells and $2000 for every new car he sells, how much would he make if he sold "u" used cars and "n" new cars? Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it.

OpenStudy (anonymous):

better

OpenStudy (anonymous):

lol, better.

OpenStudy (anonymous):

now can you help me

OpenStudy (anonymous):

As an example, consider the following. If I sell you candy bars at 2 dollars each, how much do I make if I sell you 3? Well, it would be 2*3 = 6, right? How much if I sell you 10? 2*10 = 20 If I sell you 15? 2*15 = 30 Whatever number of candy bars I sell you, you multiply that by the cost to find out how much I made. So if we make a variable b = number of candy bars then b*2 will be how much I make.

OpenStudy (anonymous):

Do you understand that? Think about how that helps you solve your problem.

OpenStudy (anonymous):

u+r=3000

OpenStudy (anonymous):

How much does he make from used cars? What if he sells 2 used cars? 5 used cars? u used cars?

OpenStudy (anonymous):

what ?